OpenStudy (anonymous):

integral of 2-3sin2x / cos2x dx?

OpenStudy (lalaly):

\[\int\limits{\frac{2-3\sin(2x)}{\cos(2x)}}dx\]??

OpenStudy (anonymous):

yup. :)

OpenStudy (lgbasallote):

have you tried separating the terms?

OpenStudy (mimi_x3):

\[\int\limits\frac{2}{\cos2x} -\frac{3\sin2x}{\cos2x} dx => \int\limits2*\frac{1}{\cos2x} - 3\tan2x dx => \int\limits2\sec2x -3\tan2x\]

OpenStudy (mimi_x3):

\[\int\limits\frac{2}{\cos2x} -\frac{3\sin2x}{\cos2x} dx => \int\limits2*\frac{1}{\cos2x} - 3\tan2x dx =>\] \[ \int\limits2\sec2x -3\tan2x\]

OpenStudy (mimi_x3):

\[\int\limits2\sec2x dx - \int\limits3\tan2x dx\]

OpenStudy (mimi_x3):

integrate term by term

OpenStudy (anonymous):

the answer is \[\ln |(1+\sin2x)|+1/2\ln|\cos2x|+C\] can you show the solution?

OpenStudy (mimi_x3):

Are you able to integrate the first one?

OpenStudy (anonymous):

nope @Mimi_x3

OpenStudy (anonymous):

can you integrate that term @Mimi_x3. will you please show me the integration. :)

OpenStudy (lalaly):

\[\int\limits{\frac{2}{\cos(2x)}}dx=\int\limits{2\sec(2x)}dx\]let u=2x du=2dx \[\int\limits{\sec(u)du}\]\[=\int\limits{secu \times \frac{secu+tanu}{secu+tanu}du}=\int\limits{\frac{secu \tan u + \sec^2u}{secu+tanu}}du\]let V=secu+tanu dv=secutanu+sec^2u so the integral becomes \[\int\limits{\frac{dv}{v}}=\ln(v)=\ln(secu+tanu)\] \[=\ln(\sec(2x)+\tan(2x))\] for the second integral \[\int\limits{\frac{\sin(2x)}{\cos(2x)}dx}\]let u=cos(2x) du=-2sin(2x)dx\[\int\limits{\frac{-1}{2}\frac{du}{u}}=-\frac{1}{2}\ln(u)=-\frac{1}{2}\ln(\cos(2x)\]

OpenStudy (mimi_x3):

im too slow :(

OpenStudy (anonymous):

its ok @Mimi_x3 thanks for your effort. :)

OpenStudy (anonymous):

also thanks for @lalaly . :)

OpenStudy (anonymous):

entonces cual es el resultado?¿

OpenStudy (anonymous):

solo veo una parte del resultado

OpenStudy (anonymous):

int (2-3sin2x)/cos 2x